What is 5 1/3% in simplest form?
Answer by Theo(13342)

this reference shows you the general procedure.
https://www.mathsisfun.com/converting-percents-fractions.html

your problem is to convert 5 and 1/3 percent to its simplest form.

to convert the percent to decimal form, divide it by 100, but keep the decimal equivalent in fraction form.

you will get 5 and 1/3 percent equals (5 + 1/3) / 100 which is equal to 5/100 + (1/3) / 100 which is equal to 5/100 + 1/300.

place both fractions under a common denominator.

you will get 5/100 + 1/300 = 15/300 + 1/300 = 16/300.

simplify the fraction to get 16/300 = 8/150 = 4/75.

i don't think it can be made simpler than that.

to see is you did it right, convert that fraction into a percent.

use your calculator to divide 4 by 75 and you will get .05333333333........

multiply that by 100 to get 5.3333333333.......%

.3333333333.......... is equal to 1/3.

your percent is therefore 5 and 1/3 %

repeating decimals can be converted to fractions by the following procedure.
http://www.basic-mathematics.com/converting-repeating-decimals-to-fractions.html

you could also have converted the fraction to percent in improper fraction form by simply multiplying it by 100.

you would have gotten 4/75 = 400/75 percent.

that percent in improper fraction form can be turned into a percent in mixed fraction form by dividing 400 by 75 to get 5 with a remainder of 25.

your percent in mixed fraction form would be 5 and 25/75 percent.

the proper fractional part of 25/75 can be reduced to 1/3.

your percent in mixed fraction form is 5 and 1/3 %.